### v1

# Matters arising: Tunnelling photons pose no challenge to Bohmian machanics

Matters arising: Tunnelling photons pose no challenge to Bohmian machanics
Yun-Fei Wang    Xiao-Yu Wang    Hui Wang Hefei National Research Center for Physical Sciences at the Microscale and School of Physical Sciences, University of Science and Technology of China, Hefei 230026, China Shanghai Research Center for Quantum Science and CAS Center for Excellence in Quantum Information and Quantum Physics, University of Science and Technology of China, Shanghai 201315, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China    Chao-Yang Lu Hefei National Research Center for Physical Sciences at the Microscale and School of Physical Sciences, University of Science and Technology of China, Hefei 230026, China Shanghai Research Center for Quantum Science and CAS Center for Excellence in Quantum Information and Quantum Physics, University of Science and Technology of China, Shanghai 201315, China Hefei National Laboratory, University of Science and Technology of China, Hefei 230088, China New Cornerstone Science Laboratory, Shenzhen 518054, China
(July 27, 2025)
Bohmian mechanics (BM) is an alternative interpretation of quantum mechanics, firstly proposed by de Broglie in 1927 [1] and further developed by Bohm in 1952 [2]. It is widely believed that BM has the same prediction on experimental outcomes with textbook Copenhagen interpretation because they obey the same fundamental Schrödinger equation in the non-relativistic regime.
However, a recent experiment by Sharoglazova et al. [3] reports an apparent discrepancy between measured energy-velocity relationships and Bohmian predictions in specific regimes. The authors contend this mismatch “casts doubt on whether the guiding equation accurately captures the temporal dynamics of the scattering process”.
In their experiment, the authors employ a high-finesse microcavity containing a dye medium. Within this cavity, the engineered height difference is implemented for the purpose of measuring speed of photons. First, a linear ramp in the main waveguide is designed to control the initial kinetic energy of photons by changing the position along this ramp at which the dye is optically pumped. Subsequently, the emitted photons impinge a constant step potential $V_{0}$, where another auxiliary waveguide starts. This auxiliary waveguide evanescently couples to the main waveguide with a fixed coupling strength $J_{0}$. This system is modeled by two coupled Schrödinger equations
$i\hbar\frac{\partial}{\partial t}\psi_{m}$ $=-\frac{\hbar^{2}}{m}\frac{d^{2}\psi_{m}}{dx^{2}}+V_{0}\psi_{m}+\hbar J_{0}(\psi_{a}-\psi_{m}),$ (1)
$i\hbar\frac{\partial}{\partial t}\psi_{a}$ $=-\frac{\hbar^{2}}{m}\frac{d^{2}\psi_{a}}{dx^{2}}+V_{0}\psi_{a}+\hbar J_{0}(\psi_{m}-\psi_{a}),$
where $m$ is the effective mass of photons considering the longitudinal resonance energy $E_{z}$ using Einstein energy-mass equivalence $E_{z}=m\tilde{c}^{2}$ ($\tilde{c}$ denotes the photon speed in the medium), $\psi_{m}$ and $\psi_{a}$ are wavefunctions in the main and auxiliary waveguides, respectively.
Figure 1: Quantum tunnelling between two coupled wavegiudes based on Schrödinger equation. a, Calculated wavefunctions $\psi_{m}$ and $\psi_{a}$ along $x$ direction. Three conditions are considered here with different $\Delta$. The upper panel shows two transmission modes, the middle panel presents one transmission and evanescent modes, and the lower panel displays two evanescent modes. b, Purple dots denotes and cyan line are experimental determined $v$ and fitted results in Ref. [3], respectively. Red line is our theoretical fitting to the experimental results in Ref. [3].
The authors perform two distinct ‘velocity’ measurements. The first method estimates a semi-classical speed $v$ through
$\rho_{a}=\left(\frac{J_{0}x}{v}\right)^{2},$ (2)
where $\rho_{a}=|\psi_{a}|^{2}/(|\psi_{a}|^{2}+|\psi_{m}|^{2})$. Thus $v$ can be determined by recording the populations $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ in the auxiliary and main waveguides at different $x$ positions, respectively. Noting that this is not velocity measurement since the direction of motion can not be determined. The authors say that this speed “represents the local speed of particles directly behind the potential step”, but the direction of this speed is still unclear. The second velocity is determined using guiding equation $\vec{v}_{s}=\frac{\hbar}{m}\nabla S$ of BM through “interfere quantum states in the step potential with their spatially mirrored images to detect phase gradients”. Based on the authors’ claim of $v_{s}=0$ for negative kinetic energy case, we believe that this $v_{s}$ should along the $x$ direction (see Methods).
Figure 2: Quantum tunnelling between two coupled waveguides based on Bohmian mechanics. Four group of velocities at $x$=5 $\mu$m (a), 10 $\mu$m (b), 20 $\mu$m (c) and 40 $\mu$m (d) are displayed. If $\Delta<-\hbar J_{0}$, both $v_{m}$ and $v_{a}$ along $x$ are zero. In the regime of $|\Delta|<\hbar J_{0}$, $v_{m}$ and $v_{a}$ are distinguished because of the different quantum potential. For $\Delta>\hbar J_{0}$, $v_{m}$ and $v_{a}$ are identical.
The discrepancy of these two measured speeds happens when $\Delta<-\hbar J_{0}$, where $\Delta=E-V_{0}+\hbar J_{0}$. The observed $v$ is nonzero and is proportional to $\sqrt{\frac{2|\Delta|}{m}}$, while measured $v_{s}$ is very close to zero. Therefore, the authors claim that this experiment raises concern on the correctness of Bohmian mechanics.
Here we reanalyze the physical process in this experiment and demonstrate that it poses no challenge to BM. First, the two measured speeds represent fundamentally distinct physical quantities; comparing them is physically unjustified and cannot serve as evidence favoring the Copenhagen interpretation over BM. Second, we theoretically prove that both the widely-held Copenhagen interpretation and BM yield identical predictions for photon tunnelling in coupled waveguides, thereby resolving the contradiction raised by Ref. [3].
We solve the equation 1 to obtain the solution
$\psi_{m}$ $\propto(e^{ik_{+}x}+e^{ik_{-}x})$ (3)
$\psi_{a}$ $\propto(e^{ik_{+}x}-e^{ik_{-}x}),$
assuming no back-propagating modes and symmetric distribution of $k_{+}$ and $k_{-}$ across both waveguides (for simplicity), where $k_{\pm}=\sqrt{\frac{2m(\Delta\mp\hbar J_{0})}{\hbar^{2}}}$. The dynamical regimes are characterized by $\Delta$. When $\Delta>\hbar J_{0}$, both $k_{+}$ and $k_{-}$ are real values, indicating there are two transmission modes in waveguide. When $|\Delta|<\hbar J_{0}$, $k_{+}$ is imaginary, representing an evanescent mode, while $k_{-}$ is real which still a transmission mode. When $\Delta<-\hbar J_{0}$, only evanescent modes exist in the waveguide, thus photon population exponentially decays. Numerical results in Fig. 1a confirm these predictions and align with experimental observations from Ref. [3].
Note that our solution differs from Ref. [3] and better describes their experimental results. For sufficiently small $x$, a Taylor expansion of $\psi_{m}$ and $\psi_{a}$ yields
$\rho_{a}=$ $\frac{2m|c_{0}|^{2}}{\hbar^{2}}(|\Delta+\hbar J_{0}|-2\sqrt{\text{Max}\{\Delta^{2}-(\hbar J_{0})^{2},0\}}$ (4)
$+|\Delta-\hbar J_{0}|)^{2}x^{2}+O(x^{3}).$
Similar to Ref. [3], $\rho_{a}$ is quadratic in $x$ but with a distinct coefficient. Fitting their defined speed $v$ (Fig. 1b) shows that our result (red line) exactly reproduces the experimental data, especially in the region $|\Delta|<\hbar J_{0}$ where a plateau appears. However, the results in Ref. [3] fail to describe this plateau.
The authors in Ref. [3] model $\rho_{a}$ as $\rho_{a}=(J_{0}x/v)^{2}$. Though they emphasize this speed $v$ “represent the the local speed of particles directly behind the potential step”, we still have concerns whether this is acceptable. Even if accepted, $v$ cannot be directly compared to the Bohmian velocity ($v_{s}$) defined by the guiding equation. First, the direction of $v$ is ambiguous—does it include a $y$-component? In contrast, $v_{s}$ is strictly along the $x$-direction (see Methods). Second, $v_{s}$ inherently satisfies the continuity equation derived from the Schrödinger equation, whereas $v$ does not. Any speed rigorously defined within the Schrödinger formalism must be equivalent to $v_{s}$ to ensure consistency with probability conservation.
Next, we derive the Bohmian velocity $v_{s}$ from continuity equation: $v_{s}(x)=\frac{\hbar}{m}\text{Im}\frac{\psi^{*}\partial_{x}\psi}{\psi^{*}\psi}$. Using the solution in equation 3, $v_{s,m}(x)$ and $v_{s,a}(x)$ can be determined. Fig. 2 shows these calculated velocities at four position ($x=5,10,20,40$ $\mu$m). In the region of $\Delta<-\hbar J_{0}$, all velocities vanish. This is expected: the wavefunction decays exponentially (purely evanescent), indicating photons are stationary along $x$. Crucially, this does not preclude tunnelling—finite probabilities $|\psi_{a}|^{2}$ and $|\psi_{m}|^{2}$ persist in the waveguides, enabling tunnelling despite zero net flow.
To definitively demonstrate this experiment poses no challenge to BM, we prove BM yields identical photon populations $\rho_{a}$ to the Schrödinger solution (equation 4; full derivation in Methods). First, substituting the wavefunction in equations 15 into the Schrödinger equations yields new quantum Hamilton-Jacobi equations 16 and continuity equations 17. The continuity equations contain a tunnelling current term $j_{0}=2J_{0}R_{a}R_{m}\text{sin}(S_{a}-S_{m})$, which accounts for inter-waveguide quantum tunnelling while preserving probability conservation. Solving the reduced continuity equation $\partial_{x}(\rho_{a,B}v_{a,x})+j_{0}=0$ with BM-derived $v_{a,x}$ and $j_{0}$ gives $\rho_{a,B}$, which is presented in equation 22 in Methods. Subsequent analysis confirms exact agreement with the Schrödinger-predicted $\rho_{a}$ (equation 4). This result indicates that (i) the Bohmian velocity $v_{a,x}$ generates the same $\rho_{a}$ as standard Schrödinger equation, and (ii) both interpretations predict identical tunnelling dynamics, refuting claims in Ref. [3].
In conclusion, the experiment in Ref. [3] poses no challenge to BM. The mistakes in Ref. [3] primarily stem from misinterpreting the semi-classical speed $v$ (equation 2) and Bohmian velocity $v_{s}$ as comparable quantities. We point out that these represent fundamentally distinct physical entities, their discrepancy cannot validate claims against BM. Furthermore, solving for $\rho_{a}$ through both Schrödinger equation and continuity equation in BM yields identical results. This confirms no interpretational distinction exists between Copenhagen and BM for quantum tunneling phenomena. In other words, any velocity defined from $\rho_{a}$ must be identical in both frameworks, meanwhile satisfy continuity equation. Last but not least, $v_{s}=0$ in the region of $\Delta<-\hbar J_{0}$ is physically expected. From the perspective of BM, photons are indeed at rest in the step potential along $x$ direction, but by no means it imply that quantum tunnelling does not occur. In fact, non-vanishing probability of evanescent wavefunction still allows quantum tunnelling to come up.